calcium comcentration of 55mg?L. What is its calcium concentration expressed in parts per million
1
Expert's answer
2012-12-06T09:44:04-0500
We firstly find theamount of substance of calcium:
n(Ca) = m/M
n(Ca) = 0.055g/40 = 0.001375 mol
using the following formula we can find theamount of substance:
n = N/NA
on this basis:
N = NA x n
N = 6.02x10^23 x 0.01337 = 8.27 x 10^20 particles of Ca
C(Ca) = 8.27 x 10^20 particles/ 1 liter = 8.27x10^20 parts
C(Ca) = 8.27x10^20*100000 liters = 8.27x10^14 parts per million